OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x2 − 8x − 42.

OpenStudy (campbell_st):

take out 2 as a factor of every term \[2(x^2 -4x - 21)\] next find the factors of -21 that add to - 4 (hint the larger factor is negative) 1 and - 21 1 - 21 = 20 3 and - 7 3 - 7 = -4 the factors to use are 3 and - 7 then its \[2x^2 - 8x - 42 = 2(x-7)(x+3)\]

OpenStudy (anonymous):

thanks